Circular No. 8431

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004fw IN UGC 4633 Further to IAUC 8417, T. Puckett and M. Peoples report the discovery

of an apparent supernova (mag 18.3) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Nov. 6.36 UT — and confirmed at mag 18.0 on frames taken on Nov. 7.30. The new object is located at α = 8^{h}52^{m}14^{s}.55, δ = +52^{o}53′30′′.4 (equinox 2000.0), which is 2′′.0 east and 5′′.8 north of the center of UGC 4633. SN 2004fw is not present images taken by Puckett on 2002 Nov. 14 and 2003 Oct. 11 (limiting mag ∼ 20.0).

SUPERNOVA 2004fx IN MCG –02-14-3 Further to IAUC 8425, S. Park and W. Li report the LOSS discovery

of an apparent supernova (mag ∼ 16.9) on an unfiltered KAIT image taken on Nov. 7.44 UT. The new object was confirmed on earlier images taken on Oct. 28.43 (mag ∼ 17.2), Nov. 1.44 (mag ∼ 17.1), and Nov. 5.43 (mag ∼ 16.9). SN 2004fx is located at α = 5^{h}11^{m}11^{s}.67, δ = −9^{o}23′31′′.6 (equinox 2000.0), which is 56′′.6 east and 11′′.2 south of the nucleus of MCG –02-14-3. A KAIT image taken on Oct. 22.42 (limiting mag ∼ 19.0) showed nothing at this position.

COMET C/2004 Q2 (MACHHOLZ) D. H. Wooden, NASA Ames Research Center; D. E. Harker, University

of California, San Diego; C. E. Woodward and M. S. Kelley, University of Minnesota; S. J. Bus and A. Tokunaga, University of Hawaii; C. Magri, University of Maine; R. P. Binzel, Massachusetts Institute of Technology; and G. Lo Curto and E. Wenderoth, European Southern Observatory (ESO), report that K-, N-, and Q-band photometry of comet C/2004 Q2 was obtained on Oct. 18.64 UT using the NASA Infrared Telescope Facility (+ SpeX guider camera). The coma was extended (FWHM = 5′′, FWZI = 12′′); aperture photometry in a 3′′-diameter aperture yields K [2.2 μm] = 8.49 ± 0.1. The comet was also observed on Oct. 23.31 with the ESO 3.6-m telescope (+ TIMMI2) at La Silla, yielding the folowing flux densities in a 3′′-diameter aperture: N_{1} (8.6 μm), 0.53 ± 0.04 Jy; Q_{1} (17.75 μm), 1.50 ± 0.14 Jy. The comet was also extended in the N_{1} band (FWHM = 3′′, FWZI = 5′′). Spitzer InfraRed Spectrograph observations of the comet are planned for Nov. 13.

2004 November 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green